# CREMA: Generalizable and Efficient Video-Language Reasoning via Multimodal Modular Fusion

### v2

@@ Abstract
Despite impressive advancements in recent multimodal reasoning approaches, they are still limited in flexibility and efficiency, as these models typically process only a few fixed modality inputs and require updates to numerous parameters. This paper tackles these critical challenges and proposes CREMA, a generalizable, highly efficient, and modular modality-fusion framework that can incorporate any new modality to enhance video reasoning. We first augment multiple informative modalities (such as optical flow, 3D point cloud, audio, thermal heatmap, and touch map) from given videos without extra human annotation by leveraging sensors or existing pre-trained models. Next, we introduce a query transformer with multiple parameter-efficient modules associated with each accessible modality. It projects diverse modality features to the LLM token embedding space, allowing the model to integrate different data types for response generation. Furthermore, we propose a novel progressive multimodal fusion design supported by a lightweight fusion module and modality-sequential training strategy. It helps compress information across various assisting modalities, maintaining computational efficiency in the LLM while improving performance. We validate our method on 7 video-language reasoning tasks assisted by diverse modalities, including conventional VideoQA and Video-Audio/3D/Touch/Thermal QA, and achieve better/equivalent performance against strong multimodal LLMs, including OneLLM, BLIP-2, and SeViLA while reducing over 90% trainable parameters. We provide extensive analyses of CREMA, including the impact of each modality on reasoning domains, the design of the fusion module, and example visualizations.  [^1] Project Page: https://CREMA-VideoLLM.github.io/.
@@ 1 Introduction
We humans understand the world through various senses, such as sight, sound, touch, and heat, allowing us to understand our environment and act accordingly. This concept has inspired the field of multimodal learning that connects various perceptions, including vision-language [1, 38, 72, 52], audio-video [19, 58], and 2D-3D joint vision [39, 22, 21, 35]. In particular, recent Multimodal Large Language Models (MLLMs) [70, 38, 42, 59] have shown promising versatility in handling multiple forms of input data, such as vision, audio, and text. These models are crucial in real-world applications that require a comprehensive understanding of multiple modalities to make decisions in various contexts. For example, autonomous vehicles rely on visual road signs, sirens, and LIDAR for navigation and safe driving. Embodied AI takes visual, heat, and touch information to complete household tasks. Similarly, educational AI enhances the learning experience by integrating various information, such as videos, speech, and textbooks.
@@
Figure 1: Overview of the CREMA architecture & training. Left: Multimodal encoders, Q-former, and LLM are kept frozen in the process. For each modality input, we extract tokens using a corresponding modality-specific adaptation module. Then, we employ the fusion module to blend and compact the obtained tokens. In the end, the LLM can leverage modality-fusion tokens to generate responses. Right: We present a modality-sequential training and modality-adaptive early exist strategy, further boosting the training efficiency while allowing faster modality adaptation.
@@
Despite their recent advancements, deploying a generic MLLM that handles multiple diverse modalities is still very challenging in terms of cost and flexibility. For different types of inputs, MLLMs have required extremely large computational budgets to update the LLM with individual encoders for modalities. Alternatively, recent efficient MLLMs using separate projection modules [74, 56, 38, 18] provide a more efficient and flexible way for multimodal reasoning. However, as each modality module contains hundreds of millions of parameters for training, this approach is still computationally intensive, and balancing as well as fusing various types of inputs becomes even more complex and costly when more modalities are introduced. Such challenges also exist in very recent pioneering works [44, 49, 45]; these models aim to integrate more diverse sensory data for compositional understanding via partial updates to the models, yet still require notable training resources to adapt to different modalities (7B for Unified-IO 2 [45]). Moreover, they focus primarily on fixed modality pairs (like 3D-text and visual-text), limiting their adaptability to new data forms and broader applications.
@@
To overcome these limitations, we present CREMA: a generalizable and highly efficient MLLM framework designed for video-language reasoning. The proposed framework enables us to incorporate any new set of modalities, including video, depth map, optical flow, surface normals, audio, thermal heatmap, touch map, 3D point cloud, notably with very few trainable parameters (4$\sim$10M for each new modality) as compared to BLIP-2 [38] ($\sim$108M) and SeViLA [69] ($\sim$216M) but higher performance due to assistance of diverse modalities. As illustrated in Figure 1 Left, given a frozen pre-trained vision-language backbone, our approach introduces modality-adaptive modules on top of the Q-Former [38] architecture, including linear projectors, low-rank adapters [23], and learnable queries. Our parameter-efficient modular design ensures that the pre-trained backbone remains unchanged and enables updates with new modalities and more advanced LLMs in the future without complex architecture changes. To enrich the input modalities, we utilize public pre-trained models to extract features from raw videos, such as depth maps, optical flow, and surface normals.
@@
Furthermore, despite the effectiveness of our versatile video reasoning framework for multimodal data, handling multiple modalities is not always beneficial; some modalities may be redundant or irrelevant to the reasoning tasks, and optimizing with all modalities simultaneously can lead to a certain deficiency, such as suboptimal convergence of the modality inputs. Besides, the LLM receives longer input contexts, which include token embeddings from all modality queries, resulting in increased computations to produce responses. To address these remaining concerns, we propose a new modality fusion technique that effectively integrates various modality tokens through a novel self-gated attention module and performs alternative modality training within a minibatch, preventing interference in the learning of modality-specific representation. Furthermore, we present a modality-adaptive early exit strategy to bypass the training of a specific modality if it is considered as converged. This further increases the efficiency of CREMA while allowing faster adaptation and maintaining comparable performance. (Please see Figure 1 Right). In the end, we allow the model to maintain GFLOPs while achieving enhanced reasoning ability, even when the LLM processes a larger number of modalities.
@@
We validate CREMA on various video-language reasoning benchmarks assisted by diverse modalities, including VideoQA (NExT-QA [66], PerceptionTest [50]), 3D-VideoQA (SQA3D [46]), Audio-VideoQA (MUSIC-AVQA [37], VGGSound [8]), Touch-VideoQA (Touch&Go [68]), Thermal-VideoQA (Thermal-IM [60]). CREMA surpasses other strong multimodal video reasoning baselines, improving fine-tuning performance by +2.4% on MUSIC-AVQA, +2.6% on SQA3D, and +0.6% on NExT-QA while reducing by 90% of the trainable parameters. CREMA also outperforms general-purpose baselines in the zero-shot setting. We further provide comprehensive analyses of varying sets of modalities, different modality fusion strategies, benefits of adding more modality, and qualitative analysis with input/response visualizations to highlight the efficiency and effectiveness of CREMA in various video-language reasoning across diverse modality inputs. Our contributions are:
@@ 2 Related Works
Beyond conventional unimodal learning, leveraging additional modalities, such as visual and audio, in learning models is increasingly popular and has demonstrated remarkable success in solving diverse tasks [77, 43, 45, 47, 64]. Vision Language Models [24, 36, 16, 9] are the most prevalent branch of multimodal learning that combine vision and language by training on massive data to understand and generate outputs involving visual and text-based information. Audio-Language Models [11, 7, 65] have been proposed for various audio-associated language tasks, e.g., spoken question answering and speech synthesis. Also, 2D-3D Joint Vision (and Language) Models [39, 22, 21, 35] aim to combine features of both two-dimensional (2D) and three-dimensional (3D) data to interpret and analyze both modalities, allowing for a more comprehensive understanding of visual information. However, these approaches are not scalable to other tasks involving different modality inputs since they focus on handling fine-grained problems with pre-defined modalities.
@@
Very recently, several works propose integrated pipelines using more than two different data sources for general-purpose reasoning [73, 18, 38, 15, 44]. MERLOT-REVERSE [73] introduces a new training objective that learns from audio, subtitles, and video frames. Given modality-specific features extracted from the encoder/tokenizer for these inputs, the joint transformer learns to predict the masked text and audio. Therefore, incorporating new types of inputs can be challenging since it requires new pre-training steps for all other modalities. X-InstructBLIP (X-BLIP) [49] integrates various modalities into a framework using frozen LLMs, employing modality-specific Q-Formers as adapters to connect different encoders. However, this method needs to train the individual Q-Former for each modality to enable modality-aligned instruction tuning, which is still resource-intensive. OneLLM [18] presents a universal encoder and projection module to align various modalities with language, its flexibility is limited in adapting new modalities, as the pre-trained projection may be impaired with unseen input format. In contrast, our CREMA method adopts an efficient and modular approach, using parameter-efficient adapters for each modality, and enhances flexibility in combining any new modalities.
@@ 3.1 Preliminaries: Q-Former
To connect various types of sensory inputs with the LLM, we adopt the Q-Former architecture originally proposed in BLIP-2 [38], a transformer [63] module that bridges the modality encoder and the LLM, similar to Perceiver [25]. It receives modality features $\mathbb{Z}$ from the encoder along with learnable queries $\bm{v}$ and produces fixed-length tokens $\bm{q}$ as output. This design enables the Q-Former to extract the most informative features from the input modality and remove any irrelevant information. It further projects obtained tokens $\bm{q}$ into the LLM’s embedding space via a fully connected layer to make them compatible. In the end, $\bm{q}$ serves as soft visual prompts [28] for the LLM. CREMA method adopts several lightweight form-adaptive modules on top of the Q-Former to integrate knowledge from different data types (e.g., video frames, audio, 3D point cloud, touch map etc.) efficiently.
@@ 3.2 Multimodal Compositional Video Reasoning via Efficient Modular Adaptation and Fusion
Our proposed method, CREMA, illustrated in Figure 1, aims to generate responses using both language (i.e., questions) and various multimodal inputs (e.g., images, audio, depth data), denoted as $\mathbb{M}=[M_{1},M_{2},...,M_{n}]$, where $n$ represents the number of accessible modalities. Throughout this paper, our CREMA method handles six modalities: video RGB frames, audio, 3D point cloud, optical flow, surface normals, and depth map, in total (up to four different modalities at once), but we note that our approach is able to process a larger number of different data types if needed. We first encode input data for each modality using the corresponding encoder. Here, we adopt several classes of publicly available pre-trained encoders for modalities, which are kept frozen, to improve training efficiency and learning generalization. Universal encoder [15, 18] can be employed as well, but it lacks the flexibility of other new modalities that are not pre-trained before. Next, we add a fully connected layer (dashed box in Figure 1) for each data type when dimension misalignment happens. It maps different modality representations into a unified feature space while avoiding incompatibilities between varying encoder architectures. We then obtain a set of multimodal features and in the next step, the Q-Former will extract informative features in $\mathbb{Z}$ into the query embeddings. More details are discussed in Section 4.1 and Section A.2.
@@
Previous studies [49, 70, 74, 20] have shown the capability of the Q-Former architecture to integrate various modalities with LLMs. However, they necessitate the individual Q-former for each modality, leading to significant parameter demands. While Q-Formers are moderately scaled at $\sim$188 million parameters, which is less than LLMs, this size becomes substantial with increasing modalities. For example, processing six different modalities would require about 1B parameters, highlighting the cost and complexity of scaling the modality with additional Q-Formers.
Hence, to deploy a lightweight, universal module capable of integrating various sensory representations, we introduce Multimodal Q-Former. This architecture integrates a Modality-specific Multi-Query Adapter (MMQA) for each modality. As illustrated in the middle of Figure 1, MMQA consists of Low-Rank Adaptation (LoRA) [23] modules [^2] We implement LoRA modules at the query and value linear projections for each self-attention layer., learnable queries, and linear projections. The intuitive design of our approach enables efficient and flexible adaptation to any specific modalities. Let $\mathbb{Z}_{m}$ be the features extracted from the data $M_{m}$ of the $m^{th}$ modality. The multimodal Q-Former propagates $\mathbb{Z}_{m}$ with the corresponding MMQA module to capture the relevant information, producing query embeddings $\bm{q}_{m}$. Given the learnable input queries $\bm{v}_{m}=\bm{v}_{m}^{0}$, we compute a linear projection at layer $i$ containing the modality-specific LoRA as follows:
@@
Our approach, which concatenates modality-adaptive queries from lightweight MMQA modules, efficiently manages multimodal reasoning tasks. However, the LLM faces increased training/inference time and computational costs due to extra input tokens proportional to the number of modalities. To prevent the query token size from growing linearly with each new modality, we introduce a novel self-gated multimodal query fusion module. We define the token embedding of video queries $\bm{q}_{V}=\bm{q}_{1}$ as a major and others $\bm{q}_{\setminus V}=\{\bm{q}_{i}\}^{n}_{i=2}$ to be supportive ones. Next, we merge supportive query embeddings through a linear projection layer $\pi(\cdot;\bm{\theta})$ to match the dimension with $\bm{q}_{V}$. Motivated by prior works [53, 30], we then perform attention on the merged query embeddings $\bar{\bm{q}}_{\setminus V}$ via self-gated operation and fuse them with $\bm{q}_{V}$:

### v4

@@ Abstract
Despite impressive advancements in recent multimodal reasoning approaches, they are still limited in flexibility and efficiency, as these models typically process only a few fixed modality inputs and require updates to numerous parameters. This paper tackles these critical challenges and proposes CREMA, a generalizable, highly efficient, and modular modality-fusion framework that can incorporate many new modalities to enhance video reasoning. We first augment multiple informative modalities (such as optical flow, 3D point cloud, audio, thermal heatmap, and touch map) from given videos without extra human annotation by leveraging sensors or existing pre-trained models. Next, we introduce a query transformer with multiple parameter-efficient modules associated with each accessible modality. It projects diverse modality features to the LLM token embedding space, allowing the model to integrate different data types for response generation. Furthermore, we propose a novel progressive multimodal fusion design supported by a lightweight fusion module and modality-sequential training strategy. It helps compress information across various assisting modalities, maintaining computational efficiency in the LLM while improving performance. We validate our method on 7 video-language reasoning tasks assisted by diverse modalities, including conventional VideoQA and Video-Audio/3D/Touch/Thermal QA, and achieve better/equivalent performance against strong multimodal LLMs, including OneLLM, BLIP-2, and SeViLA while reducing over 90% trainable parameters. We provide extensive analyses of CREMA, including the impact of each modality on reasoning domains, the design of the fusion module, and example visualizations.  [^1] Project Page: https://CREMA-VideoLLM.github.io/.
@@ 1 Introduction
We humans understand the world through various senses, such as sight, sound, touch, and heat, allowing us to understand our environment and act accordingly. This concept has inspired the field of multimodal learning that connects various perceptions, including vision-language (Alayrac et al., 2022, Li et al., 2023b, Zang et al., 2023, Radford et al., 2021), audio-video (Han et al., 2020, Tang et al., 2022), and 2D-3D joint vision (Li et al., 2020, Hou et al., 2021; 2023, Lei et al., 2024). In particular, recent Multimodal Large Language Models (MLLMs) (Yu et al., 2023b, Li et al., 2023b, Liu et al., 2023a, Tang et al., 2023a) have shown promising versatility in handling multiple forms of input data, such as vision, audio, and text. These models are crucial in real-world applications that require a comprehensive understanding of multiple modalities to make decisions in various contexts. For example, autonomous vehicles rely on visual road signs, sirens, and LIDAR for navigation and safe driving. Embodied AI takes visual, heat, and touch information to complete household tasks. Similarly, educational AI enhances the learning experience by integrating various information, such as videos, speech, and textbooks.
@@
Figure 1: Overview of the CREMA architecture & training. Left: Multimodal encoders, Q-former, and LLM are kept frozen in the process. For each modality input, we extract tokens using a corresponding modality-specific adaptation module. Then, we employ the fusion module to blend and compress the obtained multimodal tokens. In the end, the LLM uses modality-fusion tokens to generate responses. Right: We present a modality-sequential training and modality-adaptive early exist strategy, further boosting the training efficiency while allowing faster modality adaptation.
@@
Despite their recent advancements, deploying a generic MLLM that handles multiple diverse modalities is still very challenging in terms of cost and flexibility. For different types of inputs, MLLMs have required extremely large computational budgets to update the LLM with individual encoders for modalities. Alternatively, recent efficient MLLMs using separate projection modules (Zhang et al., 2023, Sun et al., 2023a, Li et al., 2023b, Han et al., 2023) provide a more efficient and flexible way for multimodal reasoning. However, as each modality module contains hundreds of millions of parameters for training, this approach is still computationally intensive, and balancing as well as fusing various types of inputs becomes even more complex and costly when more modalities are introduced. Such challenges also exist in very recent pioneering works (Liu et al., 2023c, Panagopoulou et al., 2023, Lu et al., 2023); these models aim to integrate more diverse sensory data for compositional understanding via partial updates to the models, yet still require notable training resources to adapt to different modalities (7B for Unified-IO 2 (Lu et al., 2023)). Moreover, they focus primarily on fixed modality pairs (like 3D-text and visual-text), limiting their adaptability to new data forms and broader applications. This raises the crucial research question: Can we enable MLLMs to efficiently leverage multiple modalities for Video-Language Reasoning at lower costs, similar to how humans do?.
@@
In this paper, we present CREMA: a generalizable and highly efficient MLLM framework designed for video-language reasoning with many modalities. The proposed framework enables us to incorporate any new set of modalities, including video, depth map, optical flow, surface normals, audio, thermal heatmap, touch map, 3D point cloud, notably with very few trainable parameters (4$\sim$10M for each new modality) as compared to BLIP-2 (Li et al., 2023b) ($\sim$108M) and SeViLA (Yu et al., 2023a) ($\sim$216M) but higher performance due to assistance of diverse modalities. As illustrated in Figure 1 Left, given a frozen pre-trained vision-language backbone, our approach introduces modality-adaptive modules on top of the Q-Former (Li et al., 2023b) architecture, including linear projectors, low-rank adapters (Hu et al., 2022), and learnable queries. Our parameter-efficient modular design ensures that the pre-trained backbone remains unchanged and enables updates with new modalities and more advanced LLMs in the future without complex architecture changes. To enrich the input modalities, we utilize pre-trained models to extract features from raw videos, e.g., depth maps and optical flow.
@@
Furthermore, despite the effectiveness of our versatile video reasoning framework for multimodal data, handling multiple modalities is not always beneficial; some modalities may be redundant or irrelevant to the reasoning tasks, and optimizing with all modalities simultaneously can lead to a certain deficiency, such as suboptimal convergence of the modality inputs. Besides, the LLM receives longer input contexts, which include token embeddings from all modality queries, resulting in increased computations to produce responses. To address these remaining concerns, we propose a new modality fusion technique that effectively integrates various modality tokens through a novel self-gated attention module and performs alternative modality training within a minibatch, preventing interference in the learning of modality-specific representation. Furthermore, we present a modality-adaptive early exit strategy to bypass the training of a specific modality if it is considered as converged. This further increases the efficiency of CREMA while allowing faster adaptation and maintaining comparable performance. (Please see Figure 1 Right). In the end, we allow the model to maintain GFLOPs while achieving enhanced reasoning ability, even when the LLM processes many modalities.
@@
We validate CREMA on various video-language reasoning benchmarks assisted by diverse modalities, including VideoQA (NExT-QA (Xiao et al., 2021), PerceptionTest (Pătrăucean et al., 2023)), 3D-VideoQA (SQA3D (Ma et al., 2023)), Audio-VideoQA (MUSIC-AVQA (Li et al., 2022), VGGSound (Chen et al., 2020)), Touch-VideoQA (Touch&Go (Yang et al., 2022)), Thermal-VideoQA (Thermal-IM (Tang et al., 2023b)). CREMA surpasses other strong multimodal video reasoning baselines, improving fine-tuning performance by +2.4% on MUSIC-AVQA, +2.6% on SQA3D, and +0.6% on NExT-QA while reducing by 90% of the trainable parameters. CREMA also outperforms general-purpose baselines in the zero-shot setting. We further provide comprehensive analyses of applying CREMA with different LLM, varying sets of modalities, different modality fusion strategies, benefits of adding more modality, and qualitative analysis with input/response visualizations to highlight the efficiency/effectiveness of CREMA in various video-language reasoning across diverse modalities. Our contributions are summarized as follows:
@@
  • With CREMA framework, we conduct extensive studies to provide a comprehensive analysis of how multimodal information helps video-centric language reasoning tasks.
@@ 2 Related Works
Beyond conventional unimodal learning, leveraging additional modalities, such as visual and audio, in learning models is increasingly popular and has demonstrated remarkable success in solving diverse tasks (Zhu et al., 2024, Liu et al., 2023b, Lu et al., 2023, Moon et al., 2023, Wang et al., 2024a). Vision Language Models (Huang et al., 2023, Li et al., 2023a, Gong et al., 2023, Chen et al., 2023b) are the most prevalent branch of multimodal learning that combine vision and language by training on massive data to understand and generate outputs involving visual and text-based information. Audio-Language Models (Chuang et al., 2020, Castellon et al., 2021, Wang et al., 2023) have been proposed for various audio-associated language tasks, e.g., spoken question answering and speech synthesis. Also, 2D-3D Joint Vision (and Language) Models (Li et al., 2020, Hou et al., 2021; 2023, Lei et al., 2024) aim to combine features of both two-dimensional (2D) and three-dimensional (3D) data to interpret and analyze both modalities, allowing for a more comprehensive understanding of visual information. However, these approaches are not scalable to other tasks involving different modality inputs since they focus on handling fine-grained problems with pre-defined modalities.
@@
Very recently, several works propose integrated pipelines using more than two different data sources for general-purpose reasoning (Zellers et al., 2022, Han et al., 2023, Li et al., 2023b, Girdhar et al., 2023, Liu et al., 2023c). MERLOT-REVERSE (Zellers et al., 2022) introduces a new training objective that learns from audio, subtitles, and video frames. Given modality-specific features extracted from the encoder/tokenizer for these inputs, the joint transformer learns to predict the masked text and audio. Therefore, incorporating new types of inputs can be challenging since it requires new pre-training steps for all other modalities. X-InstructBLIP (X-BLIP) (Panagopoulou et al., 2023) integrates various modalities into a framework using frozen LLMs, employing modality-specific Q-Formers as adapters to connect different encoders. However, this method needs to train the individual Q-Former for each modality to enable modality-aligned instruction tuning, which is still resource-intensive. MultiPLY (Hong et al., 2024) is a multisensory embodied LLM for interaction within 3D environments with fixed modality set. X-VILA (Ye et al., 2024) is an omni-modality model aimed at cross-modality alignment, understanding, and generation. It concentrates on large-scale cross-modality alignment. OneLLM (Han et al., 2023) presents a universal encoder and projection module to align various modalities with language, its flexibility is limited in adapting new modalities, as the pre-trained projection may be impaired with unseen input format. Our CREMA method adopts an efficient and modular approach, using parameter-efficient adapters for each modality, and enhances flexibility in combining any new modalities.
@@ 3.1 Preliminaries: Q-Former
To connect various types of sensory inputs with the LLM, we adopt the Q-Former architecture originally proposed in BLIP-2 (Li et al., 2023b), a transformer (Vaswani et al., 2017) module that bridges the modality encoder and the LLM, similar to Perceiver (Jaegle et al., 2021). It receives modality features $\mathbb{Z}$ from the encoder along with learnable queries $\bm{v}$ and produces fixed-length tokens $\bm{q}$ as output. This design enables the Q-Former to compress image tokens into a fixed-length set of query tokens, facilitating efficient processing of video inputs while preserving critical holistic information. It further projects obtained tokens $\bm{q}$ into the LLM’s embedding space via a fully connected layer to make them compatible. In the end, $\bm{q}$ serves as soft visual prompts (Jia et al., 2022) for the LLM. CREMA method adopts several lightweight form-adaptive modules on top of the Q-Former to integrate knowledge from different data types (e.g., video frames, audio, 3D point cloud, touch map etc.) efficiently.
@@ 3.2 Multimodal Compositional Video Reasoning via Efficient Modular Adaptation and Fusion
Our proposed method, CREMA, illustrated in Figure 1, aims to generate responses using both language (i.e., questions) and various multimodal inputs (e.g., images, audio, depth data), denoted as $\mathbb{M}=[M_{1},M_{2},...,M_{n}]$, where $n$ represents the number of accessible modalities. Throughout this paper, our CREMA method handles six modalities: video RGB frames, audio, 3D point cloud, optical flow, surface normals, and depth map, in total (up to four different modalities at once), but we note that our approach is able to process a larger number of different data types if needed. We first encode input data for each modality using the corresponding encoder. Here, we adopt several classes of publicly available pre-trained encoders for modalities, which are kept frozen, to improve training efficiency and learning generalization. Universal encoder (Girdhar et al., 2023, Han et al., 2023) can be employed as well, but it lacks the flexibility of other new modalities that are not pre-trained before. Next, we add a fully connected layer (dashed box in Figure 1) for each data type when dimension misalignment happens. It maps different modality representations into a unified feature space while avoiding incompatibilities between varying encoder architectures. We then obtain a set of multimodal features and in the next step, the Q-Former will extract informative features in $\mathbb{Z}$ into the query embeddings. More details are discussed in Section 4.1 and Section A.2.
@@
Previous studies (Panagopoulou et al., 2023, Yu et al., 2023b, Zhang et al., 2023, Hong et al., 2023) have shown the capability of the Q-Former architecture to integrate various modalities with LLMs. However, they necessitate the individual Q-former for each modality, leading to significant parameter demands. While Q-Formers are moderately scaled at $\sim$188 million parameters, which is less than LLMs, this size becomes substantial with increasing modalities. For example, processing six different modalities would require about 1B parameters, highlighting the cost and complexity of scaling the modality with additional Q-Formers.
Hence, to deploy a lightweight, universal module capable of integrating various sensory representations, we introduce Multimodal Q-Former. This architecture integrates a Modality-specific Multi-Query Adapter (MMQA) for each modality. As illustrated in the middle of Figure 1, MMQA consists of Low-Rank Adaptation (LoRA) (Hu et al., 2022) modules [^2] We implement LoRA modules at the query and value linear projections for each self-attention layer., learnable queries, and linear projections. The intuitive design of our approach enables efficient and flexible adaptation to any specific modalities. Let $\mathbb{Z}_{m}$ be the features extracted from the data $M_{m}$ of the $m^{th}$ modality. The multimodal Q-Former propagates $\mathbb{Z}_{m}$ with the corresponding MMQA module to capture the relevant information, producing query embeddings $\bm{q}_{m}$. Given the learnable input queries $\bm{v}_{m}=\bm{v}_{m}^{0}$, we compute a linear projection at layer $i$ containing the modality-specific LoRA as follows:
@@
Our approach, which concatenates modality-adaptive queries from lightweight MMQA modules, efficiently manages multimodal reasoning tasks. However, the LLM faces increased training/inference time and computational costs due to extra input tokens proportional to the number of modalities. To prevent the query token size from growing linearly with each new modality, we introduce a novel self-gated multimodal query fusion module (Assistance Modality Fusion in Figure 1). We define the token embedding of video queries $\bm{q}_{V}=\bm{q}_{1}$ as a major and others $\bm{q}_{\setminus V}=\{\bm{q}_{i}\}^{n}_{i=2}$ to be supportive ones. Next, we merge supportive query embeddings through a linear projection layer $\pi(\cdot;\bm{\theta})$ to match the dimension with $\bm{q}_{V}$. Motivated by Ramachandran et al. (2017), we then perform attention on the merged query embeddings $\bar{\bm{q}}_{\setminus V}$ via self-gated operation and fuse them with $\bm{q}_{V}$: